The behavior of real gas is usually depicted by plotting compressibility factor z versus p at constant temperature. At high temperature and high pressure z is usually more than one(z>1). This fact can be explained by vanderwaals equation when
a)constant a is negligible and not b
b)constant b is not negligible and not a
c)both constant a and b are negligible
d)both constant a and b are not negligible

Grade:11

1 Answers

vikas askiitian expert
13 years ago

at high pressure , molar volume is low so we cannot neglect b ...

at high pressure , correction term in pressure (a/v2) can be neglected so

(p+a/v2)(v-b) = RT

this equation becomes

P(v-b) = RT

PV - Pb = RT

PV/RT - Pb/RT = 1

Z = 1 + Pb/RT

or

Z > 1

option a) is correct